### v1

@@ 3.2.1 RQ-A1 Approach
3.2.1.1 Match code changes
The goal of this step is to compare the code changes the LLM made (Changellm) with the developer ground truth changes (Changedev), as well as manually finding alternative correct changes.
Auto change matching: We first attempt to automatically match as many changes as possible while ensuring precision i.e., not falsely marking a Changellm as correct. Therefore, we consider only exact syntactic matches between Changedev and Changellm, ignoring formatting differences. We find exact matches by identifying the AST nodes related to Changedev in Filedev and then matching them to corresponding nodes in Filellm. If there are potentially multiple matches, we check the containing function and use proximity based heuristics demonstrated below. We ensure that the matched node sets translate to the same code string using ast.unparse docast .
Consider Changedev D1 in 1(c) which uses ClientSession() and get() from aiohttp. In Filellm, these APIs appear in multiple locations, but only Lines 15, 16, 21, and 22 are in the same containing function fetch_flights as Changedev. Based on line number proximity, we identify two node sets L2 and L4 as potential matches for this Changedev. By comparing the code strings, we find that L2 exactly matches Changedev. By the end of the auto change matching step, we get a set of matched pairs of Changedev and Changellm, a set of unmatched Changedev (uChangedev), and a set of unmatched Changellm (uChangellm).
Manual review: The manual review step focuses on reviewing uChangedev and uChangellm. We use a three-way diff viewer to manually compare Filepre, Filedev, and Filellm for each file with at least one uChangedev or uChangellm. We use the online API documentation as well as the source code of the source and target libraries to understand the correct API usage in the context of the file. For 1(c), we manually find that L4 is a correct alternative to D2, noting that the LLM omitted the default value (True) of the allow_redirects argument, unlike the developer. L8 is a correct replacement of D3, where the LLM performed a refactoring. However, we find that the LLM used the function get instead of post at line 33, making L0 an incorrect alternative of D4.
@@ 3.2.1.1 Match code changes
We perform a multi-reviewer manual review to ensure the quality of the matching process. The thesis author first manually reviews 190 code changes from one model. Two collaborators then review about half of those code changes each, not knowing the thesis author’s decisions. Therefore, each code change is reviewed by two reviewers independently. We then measure the agreement using Cohen’s Kappa cohen1960coefficient to quantify the level of agreement between the reviewers. We do not reach substantial agreement, so we discuss and resolve disagreements and update our coding guideline. The two pairs of reviewers then independently review another 111 code changes, achieving a Cohen’s Kappa score of 0.83 (almost perfect) and 0.73 (substantial)  landis1977measurement . Accordingly, we proceed with only the thesis author reviewing the remaining changes.
3.2.1.2 Determine migration status
Based on the matching of Changedev and Changellm, we determine each code change status as follows (example changes refer to 1(c)).
@@ 3.2.1.2 Determine migration status
3.2.2 RQ-A1 Findings: migration level correctness
@@
RQ-A1, Migration Level: All three LLMs were able to perform correct migrations with 4o performing best: 94% of its migrations were partially correct and 57% were fully correct.
3.2.3 RQ-A1 Findings: code change level correctness
Table 2: Correctness of PyMigBench migrations (RQ1, migration level)
@@ 3.2.3 RQ-A1 Findings: code change level correctness
We now present the correctness results at the individual code change level. The 314 migrations we use for our evaluation contain a total of 2,910 code changes. However, when there is a response failure or syntax error in the LLM’s migration, we cannot evaluate the correctness of the corresponding code changes in that file. Accordingly, in this analysis level, we exclude 1,999 code changes from LLama, 944 from Mini, and 914 from 4o. This leaves us with 911, 1,966, and 1,996 code changes to evaluate for the three LLMs, respectively.
3.2.3.1 Overall code change correctness
Table 3: Correctness of PyMigBench migration-related code changes (RQ1, overall code change level)
@@ 3.2.3.1 Overall code change correctness
3.2.3.2 Code change correctness by category
Table 4: Correctness of code changes across PyMigTax categories (RQ-A1, code change level by category)
@@ 3.2.3.2 Code change correctness by category
PyMigTax category Frequency in PyMigBench LLama 3.1 GPT-4o mini GPT-4o Migs LPs $C$s #Eval %Cor #Eval %Cor #Eval %Cor Program elements Function call 229 (71%) 105 (78%) 1,804 (60%) 518 88% 1,093 87% 1,092 93% Import 318 (99%) 132 (99%) 732 (24%) 276 95% 580 94% 593 98% Decorator 39 (12%) 13 (10%) 411 (14%) 197 87% 335 90% 353 87% Attribute 54 (17%) 33 (25%) 183 (6.1%) 63 83% 103 78% 103 89% Type 17 (5.3%) 13 (10%) 55 (1.8%) 12 75% 34 85% 34 88% Exception 16 (5.0%) 13 (10%) 29 (1.0%) 17 65% 29 86% 29 93% Function ref 4 (1.2%) 4 (3.0%) 12 (0.4%) 3 33% 12 58% 12 100% Properties NoProps 57 (18%) 36 (27%) 262 (8.8%) 43 84% 150 88% 150 97% ElemNC 188 (59%) 88 (66%) 1,025 (34%) 398 84% 773 85% 771 90% ArgAdd 93 (29%) 49 (37%) 578 (19%) 150 85% 244 80% 244 88% ArgDel 66 (21%) 37 (28%) 349 (12%) 175 85% 234 87% 234 89% ArgTrans 81 (25%) 38 (28%) 317 (11%) 115 77% 201 76% 220 77% ParamAdd 13 (4.0%) 1 (0.7%) 127 (4.2%) 110 87% 127 95% 127 90% ArgNC 17 (5.3%) 11 (8.2%) 112 (3.7%) 20 60% 58 59% 58 84% AsyncTrans 13 (4.0%) 4 (3.0%) 50 (1.7%) 35 83% 50 90% 50 92% OutTrans 16 (5.0%) 15 (11%) 49 (1.6%) 25 88% 49 78% 49 92% Cardinality One-to-One 200 (62%) 98 (73%) 1,552 (52%) 408 86% 877 84% 894 91% One-to-Zero 50 (16%) 25 (19%) 297 (10%) 99 99% 290 99% 290 100% Zero-to-One 35 (11%) 18 (13%) 198 (6.6%) 34 97% 55 87% 55 89% Higher Card. 91 (28%) 45 (34%) 210 (7.0%) 94 70% 164 79% 164 84% One-to-Many 52 (16%) 23 (17%) 142 (4.8%) 61 69% 103 78% 103 80% Many-to-One 30 (9.3%) 24 (18%) 41 (1.4%) 16 62% 34 82% 34 91% Many-to-Many 21 (6.5%) 11 (8.2%) 27 (0.9%) 17 82% 27 78% 27 89% Total 321 134 2,989 911 89% 1,966 89% 1,996 94%
@@
To understand if some types of code changes are more difficult to migrate, Table 4 shows the distribution of code change correctness across the different PyMigTax categories. The three columns in group “Frequency in PyMigBench” shows how often each PyMigTax category appears in all migrations, library pairs and code changes in PyMigBench, which provides perspective on the category’s impact in practice. This information is the same for each LLM since it describes the existing data in PyMigBench. For example, 1,804 (60%) of all 2,989 code changes involve function calls. The next three column groups show the performance of each model. Since each LLM has a different number of evaluated code changes, we show the number of evaluated code changes from each category (#Eval) alongside the proportion of these code changes that are correct (%Cor). For example, the Function call row under the Program elements group shows that LLama’s syntactically correct migrations included 518 code changes that involve function calls, and that 88% of these code changes were correctly migrated.
@@
Program elements: Function calls are the most common program elements in terms of code changes (60%), and second most common in terms of migrations (71%) and library pairs (78%). All three LLMs correctly migrate a high proportion of these changes, with 88%, 87%, and 93% of the function calls being correctly migrated by the three LLMs. Almost all migrations and library pairs in PyMigBench require migrating import statements, a task that all three LLMs perform well, especially 4o correctly migrates nearly all of them (98%).
@@
Properties: The property no properties indicates that the source and target APIs are identical, and hence no changes are required. While 4o correctly leaves almost all of these APIs unchanged (97%), LLama and Mini incorrectly changed several of them, resulting in 84% and 88% correct, respectively.
@@
All three LLMs performed better or equally well in argument deletion compared to those with argument addition, which is expected as adding an argument requires the LLM to find a suitable replacement, while deleting an argument is a simpler task. The Argument transformation property indicates that an argument requires various changes, such as changing the type or value. The three LLMs correctly migrate only 76-77% of the argument transformation changes. We discuss specific failure instances in Section 3.4.
The property parameter addition to decorated function is found in migrations in just one library, click library, where some decorators require adding a parameter to the decorated function libclick . All three LLMs perform well in migrating these changes; interestingly, Mini performs better than 4o (95% vs. 90%). Mini also performed better than 4o in migrating decorators, which are common in the click library. These are the only two categories where 4o does not perform the best.
@@
Cardinality: One-to-One code changes are the most common in PyMigBench, and all three LLMs perform well in migrating them. One-to-Zero code changes are the ones where a source API needs to be removed, but no target needs to be added. Because it is a simple delete operation, all three LLMs migrate all or almost all of these changes correctly. Its counterpart, Zero-to-One code changes, are the ones where a target needs to be added, but no source needs to be removed. Interestingly, LLama performs the best in migrating these changes (97% correct), making this the only category where it performs better than the other models (87% and 89% correct). While previous research showed that higher-cardinality code changes are generally difficult compared to one-to-one code changes alrubaye2018automating ; wang2016transforming ; zhang2020deep ; HuangMappingAPI2024 , the PyMigBench dataset shows that higher cardinality changes are frequent with 34% of the library pairs requiring at least one higher-cardinality code change. Overall, we find that the LLMs are reasonably successful at migrating them, with 70%, 79%, and 84% correctness rates.
RQ-A1: Code Change Level LLMs correctly migrate a high proportion of most code changes, with 4o achieving 94% correct code changes. The models perform reasonably well in migrating difficult higher-cardinality code changes (4o gets 84% correct), but struggle relatively more with argument transformation and argument name change.
3.3 RQ-A2 How many migrations pass unit tests?
We now assess the same LLM PyMigBench migrations from RQ1 but using the unit tests available in the corresponding repositories. A test-based evaluation that runs the code ensures that the migration achieves the expected behavior. Note that we should not expect that an LLM migration improves the rate of passing tests; a successful migration means that each test that previously passed on codedev must still pass on codellm. We now explain our evaluation set up.
3.3.1 Preparing the code
3.3.2 Setting up the virtual environment
3.3.3 Running the tests and coverage
Once the virtual environment is ready, we run all the unit tests while measuring coverage in the repository on codedev. If the run has errors, we read the error log and try to fix the errors, and run the tests again. We maintain a configuration file where we record any project-specific configurations or commands we used. We include this information in our artifact.
We find that 218 out of the 314 migrations have at least one test. Among these, 172 migrations had unresolvable errors. The failures are primarily due to missing dependencies, specially for older projects. The remaining 46 migrations have tests that we are able to successfully run on codedev.
However, for the tests to be useful to validate migration, they must cover the migration-related code changes. Using the coverage report, we find that only 27 migrations have at least one test that covers the code changes, and among them, 25 migrations have at least one test that passes on codedev. We run the entire test suite on codellm for these 25 migrations for each model using the same randomly selected LLM run used in RQ-A1.
3.3.4 Determining migration status
  • Correct: All passing tests in codedev also pass in codellm.
3.3.5 RQ-A2 Findings
Table 5: Correctness of 25 PyMigBench migrations using their available unit tests (RQ2)
@@ 3.3.5 RQ-A2 Findings
RQ-A2 Out of 25 PyMigBench migrations with unit tests covering the migration-related code changes, the LLMs correctly migrated 36%-64%, with 4o being the highest.
3.4 Implications and discussion
3.4.1 Are LLMs suitable for library migration?
We look into some specific examples to understand the promises and limitations of LLMs for library migration.
3.4.1.1 The good
We start with the promising side of using LLMs for library migration. In addition to the high overall correctness of individual code changes and migrations in RQ1, our evaluation reveals that the evaluated LLMs can migrate code changes that were traditionally known as difficult to migrate xu2019meditor . We discuss some examples.
Different API styles: The library argparse libargparse provides API functions for parsing command line arguments, while click libclick provides decorators. Figure 2 shows a migration example between these libraries. Despite the completely different API styles, the three LLMs correctly migrate 94% of code changes between this library pair.
Figure 2: 4o correctly migrating from argparse to click, libraries using different API styles.
Higher cardinality migrations As shown in RQ2, LLama, Mini, and 4o successfully migrated 70%, 79%, and 84% of the higher cardinality code changes, respectively, which the literature always viewed as complex migrations xu2019meditor ; alrubaye2020learning . Figure 3 shows an example of a many-to-many migration from twitter to tweepy done by LLama. In addition to correctly replacing the functions Twitter(), and OAuth() with OAuthHandler(), set_access_token(), and API() with correct arguments, LLama also splits the code into multiple statements, making the code more readable.
twitter to tweepy, cloudbotirc/cloudbot/f8243223/twitter.py

### v2

@@ 3.1.2 Matching LLM migrations with developer migration
A. Auto change matching
We first attempt to automatically match as many LLM changes (Changellm) as possible with developer change (Changedev), while ensuring precision i.e., not falsely marking an incorrect Changellm as correct. Therefore, we consider only exact syntactic matches between Changedev and Changellm, ignoring formatting differences. We find exact matches by identifying the AST nodes related to Changedev in Filedev and then matching them to corresponding nodes in Filellm. If there are potentially multiple matches, we check the containing function and use proximity based heuristics demonstrated below. We ensure that the matched node sets translate to the same code string using ast.unparse.
Consider Changedev D1 in 2(c) which uses ClientSession() and get() from aiohttp. In Filellm, these APIs appear in multiple locations, but only Lines 15, 16, 21, and 22 are in the same containing function fetch_flights as Changedev. Based on line number proximity, we identify two node sets L2 and L4 as potential matches for this Changedev. By comparing the code strings, we find that L2 exactly matches Changedev. By the end of the auto change matching step, we get a set of matched pairs of Changedev and Changellm, a set of unmatched Changedev (uChangedev), and a set of unmatched Changellm (uChangellm).
B. Manually reviewing unmatched changes
The automatic matching ensures that no change is incorrectly matched, but it may miss some correct matches due to syntactic differences, or alternative correct implementations. The manual review step focuses on reviewing uChangedev and uChangellm. We use the online API documentation as well as the source code of the source and target libraries to understand the correct API usage in the context of the file. For 2(c), we manually find that L4 is a correct alternative to D2, noting that the LLM omitted the default value (True) of the allow_redirects argument, unlike the developer. L8 is a correct replacement of D3, where the LLM performed a refactoring. However, we find that the LLM used the function get instead of post at line 33, making L0 an incorrect alternative of D4.
@@ B. Manually reviewing unmatched changes
We perform a multi-reviewer manual review to ensure the quality of the matching process. The first author first manually reviews 190 code changes from one model. Two other authors then review about half of those code changes each independently. We then measure the agreement using Cohen’s Kappa Cohen (1960) to quantify the level of agreement between the reviewers. We do not reach substantial agreement, so we discuss and resolve disagreements and update our coding guideline. The two pairs of reviewers then independently review another 111 code changes, achieving a Cohen’s Kappa score of 0.83 (almost perfect) and 0.73 (substantial)  Landis and Koch (1977). Accordingly, we proceed with only the first author reviewing the remaining changes.
C. Determining code change status
Based on the matching of Changedev and Changellm, we determine each code change status as follows (example changes refer to 2(c)).
@@ D. Determining migration status
3.1.3 Evaluating LLM migrations using unit tests
In RQ-1.2, we assess the same LLM PyMigBench migrations from RQ-1.1, but using the unit tests available in the corresponding repositories. A test-based evaluation that runs the code ensures that the migration preserves the original behavior. We now explain the process below.
Once the virtual environment is ready, we run all the unit tests while measuring coverage in the repository on Codedev. If the run has errors, we read the error log and try to fix the errors, and run the tests again. Note that we only fix configuration or environment-related errors, not code errors. For example, if a dependency is missing, we install it; if the project requires a specific Python version, we set up the virtual environment with that version. We maintain a configuration file where we record any project-specific configurations or commands we used. We include this information in our artifact.
We find that 218 out of the 314 migrations have at least one test. Among these, 172 migrations had unresolvable errors. The failures are primarily due to missing dependencies, especially for older projects. The remaining 46 migrations have tests that we are able to successfully run on Codedev.
However, for the tests to be useful to validate migration, they must cover the migration-related code changes. Using the coverage report, we find that only 27 migrations have at least one test that covers the code changes, and among them, 25 migrations have at least one test that passes on Codedev. Thus, the evaluation in RQ-1.2 focuses only on these 25 migrations. We run the entire test suite on Codellm for these 25 migrations for each model using the same randomly selected LLM run used in RQ-1.1.
  • Correct: All passing tests in Codedev also pass in Codellm.
3.2 RQ-1.1 How similar are the LLM migrations to the benchmark migrations?
We now discuss the results of RQ-1.1. We first present the migration level results, then present the overall code change level results, and finally analyze the correctness of code changes by PyMigTax categories.
3.2.1 Migration level correctness
@@
RQ-1.1: Migration level All three LLMs were able to perform correct migrations with 4o performing best: 94% of its migrations were partially correct and 57% were fully correct.
3.2.2 Overall code change level correctness
Table 1: Correctness of PyMigBench migrations (RQ-1.1, migration level)
@@ 3.2.2 Overall code change level correctness
The 314 migrations we use for our evaluation contain a total of 2,910 code changes. However, when there is a response failure or syntax error in the LLM’s migration, we cannot evaluate the correctness of the corresponding code changes in that file. Accordingly, in this analysis level, we exclude 1,999 code changes from Llama, 944 from Mini, and 914 from 4o. This leaves us with 911, 1,966, and 1,996 code changes to evaluate for the three LLMs, respectively.
Table 2: Correctness of PyMigBench migration-related code changes (RQ-1.1, overall code change level)
@@
3.2.3 Code change correctness by category
Table 3: Correctness of code changes across PyMigTax categories (RQ-1.1, code change level by category)
@@ 3.2.3 Code change correctness by category
PyMigTax category Frequency in PyMigBench Llama 3.1 GPT-4o mini GPT-4o
Migs LPs $C$s #Eval %Cor #Eval %Cor #Eval %Cor
Program elements
Function call 229 (71%) 105 (78%) 1,804 (60%) 518 88% 1,093 87% 1,092 93%
Import 318 (99%) 132 (99%) 732 (24%) 276 95% 580 94% 593 98%
Decorator 39 (12%) 13 (10%) 411 (14%) 197 87% 335 90% 353 87%
Attribute 54 (17%) 33 (25%) 183 (6.1%) 63 83% 103 78% 103 89%
Type 17 (5.3%) 13 (10%) 55 (1.8%) 12 75% 34 85% 34 88%
Exception 16 (5.0%) 13 (10%) 29 (1.0%) 17 65% 29 86% 29 93%
Function ref 4 (1.2%) 4 (3.0%) 12 (0.4%) 3 33% 12 58% 12 100%
Properties
NoProps 57 (18%) 36 (27%) 262 (8.8%) 43 84% 150 88% 150 97%
ElemNC 188 (59%) 88 (66%) 1,025 (34%) 398 84% 773 85% 771 90%
ArgAdd 93 (29%) 49 (37%) 578 (19%) 150 85% 244 80% 244 88%
ArgDel 66 (21%) 37 (28%) 349 (12%) 175 85% 234 87% 234 89%
ArgTrans 81 (25%) 38 (28%) 317 (11%) 115 77% 201 76% 220 77%
ParamAdd 13 (4.0%) 1 (0.7%) 127 (4.2%) 110 87% 127 95% 127 90%
ArgNC 17 (5.3%) 11 (8.2%) 112 (3.7%) 20 60% 58 59% 58 84%
AsyncTrans 13 (4.0%) 4 (3.0%) 50 (1.7%) 35 83% 50 90% 50 92%
OutTrans 16 (5.0%) 15 (11%) 49 (1.6%) 25 88% 49 78% 49 92%
Cardinality
One-to-One 200 (62%) 98 (73%) 1,552 (52%) 408 86% 877 84% 894 91%
One-to-Zero 50 (16%) 25 (19%) 297 (10%) 99 99% 290 99% 290 100%
Zero-to-One 35 (11%) 18 (13%) 198 (6.6%) 34 97% 55 87% 55 89%
Higher Card. 91 (28%) 45 (34%) 210 (7.0%) 94 70% 164 79% 164 84%
One-to-Many 52 (16%) 23 (17%) 142 (4.8%) 61 69% 103 78% 103 80%
Many-to-One 30 (9.3%) 24 (18%) 41 (1.4%) 16 62% 34 82% 34 91%
Many-to-Many 21 (6.5%) 11 (8.2%) 27 (0.9%) 17 82% 27 78% 27 89%
Total 321 134 2,989 911 89% 1,966 89% 1,996 94%
@@
To understand if some types of code changes are more difficult to migrate, Table 3 shows the distribution of code change correctness across the different PyMigTax categories. The three columns in group “Frequency in PyMigBench” shows how often each PyMigTax category appears in all migrations (Migs), library pairs (LPs) and code changes (CCs) in PyMigBench, which provides perspective on the category’s impact in practice. This information is the same for each LLM since it describes the existing data in PyMigBench. For example, 1,804 (60%) of all 2,989 code changes involve function calls. The next three column groups show the performance of each model. Since each LLM has a different number of evaluated code changes, we show the number of evaluated code changes from each category (#Eval) alongside the proportion of these code changes that are correct (%Cor). For example, the Function call row under the Program elements group shows that Llama’s syntactically correct migrations included 518 code changes that involve function calls, and that 88% of these code changes were correctly migrated.
A. Program elements
@@
Function calls are the most common program elements in terms of code changes (60%), and second most common in terms of migrations (71%) and library pairs (78%). All three LLMs correctly migrate a high proportion of these changes, with 88%, 87%, and 93% of the function calls being correctly migrated by the three LLMs. Almost all migrations and library pairs in PyMigBench require migrating import statements, a task that all three LLMs perform well, especially 4o correctly migrates nearly all of them (98%).
@@
B. Properties
The property no properties indicates that the source and target APIs are identical, and hence no changes are required. While 4o correctly leaves almost all of these APIs unchanged (97%), Llama and Mini incorrectly changed several of them, resulting in 84% and 88% correct, respectively.
@@ B. Properties
All three LLMs performed better or equally well in argument deletion compared to those with argument addition, which is expected as adding an argument requires the LLM to find a suitable replacement, while deleting an argument is a simpler task. The Argument transformation property indicates that an argument requires various changes, such as changing the type or value. The three LLMs correctly migrate only 76-77% of the argument transformation changes.
The property parameter addition to decorated function is found in migrations in just one library, click library, where some decorators require adding a parameter to the decorated function Pallets (2024). All three LLMs perform well in migrating these changes; interestingly, Mini performs better than 4o (95% vs. 90%). Mini also performed better than 4o in migrating decorators, which are common in the click library. These are the only two categories where 4o does not perform the best.
@@
C. Cardinality
One-to-One code changes are the most common in PyMigBench, and all three LLMs perform well in migrating them. One-to-Zero code changes are the ones where a source API needs to be removed, but no target needs to be added. Because it is a simple delete operation, all three LLMs migrate all or almost all of these changes correctly. Its counterpart, Zero-to-One code changes, are the ones where a target needs to be added, but no source needs to be removed. Interestingly, Llama performs the best in migrating these changes (97% correct), making this the only category where it performs better than the other models (87% and 89% correct). While previous research showed that higher-cardinality code changes are generally difficult compared to one-to-one code changes Alrubaye and Mkaouer (2018); Wang et al. (2016); Zhang et al. (2020); Huang et al. (2024), the PyMigBench dataset shows that higher cardinality changes are frequent with 34% of the library pairs requiring at least one higher-cardinality code change. Overall, we find that the LLMs are reasonably successful at migrating them, with 70%, 79%, and 84% correctness rates.
RQ-1.1: Code change level LLMs correctly migrate a high proportion of most code changes, with 4o achieving 94% correct code changes. The models perform reasonably well in migrating difficult higher-cardinality code changes (4o gets 84% correct), but struggle relatively more with argument transformation and argument name change.
3.3 RQ-1.2 How many migrations pass unit tests?
Table 4: Correctness of 25 PyMigBench migrations using their available unit tests (RQ-1.2)
@@ C. Cardinality
RQ-1.2 Out of 25 PyMigBench migrations with unit tests covering the migration-related code changes, the LLMs correctly migrated 36%-64%, with 4o being the highest.
3.4 Strengths of LLM for library migration
Overall, the results of this empirical study highlight the high overall correctness of LLM migrations. We now discuss the key strengths we observe.
Higher cardinality code changes
As shown in RQ-1.1, Llama, Mini, and 4o successfully migrated 70%, 79%, and 84% of the higher cardinality code changes, respectively, which the literature always viewed as complex migrations Xu et al. (2019); Alrubaye et al. (2020). Figure 3 shows an example of a many-to-many migration from twitter to tweepy done by Llama. In addition to correctly replacing the functions Twitter(), and OAuth() with OAuthHandler(), set_access_token(), and API() with correct arguments, Llama also splits the code into multiple statements, making the code more readable.
Before migration (using twitter)
4tw_api = Twitter(
5 auth = OAuth(oauth_token, oauth_secret,
6 consumer_key, consumer_secret))
7
8 ÂX
Llama’s migration (using tweepy)
4auth = OAuthHandler(
5 consumer_key, consumer_secret)
6auth.set_access_token(
7 oauth_token, oauth_secret)
8tw_api = API(auth)